# Scaling Laws in Patchification: An Image Is Worth 50,176 Tokens And More

### v1

@@ Abstract
1 Johns Hopkins University 2 UC Berkeley 3 University of Florida 4 UC Santa Cruz
@@ 1 Introduction
Thanks to the rapid advancements in hardware, efficient attention mechanisms (Dao et al., 2022; Kwon et al., 2023), as well as linear-complexity structures (Katharopoulos et al., 2020; Peng et al., 2023; Gu & Dao, 2023), we can now extensively validate the impact of patchification at a standard input size (e.g., 224224 for ImageNet) with manageable computing resources. We conduct a series of straightforward scaling experiments on patchification, gradually reducing the model’s patch size from the typical 1616 down to 11 to lower the compression rate and observe how performance changes. We employ both ViT and Adventurer (Wang et al., 2024b), a Mamba-based (Dao & Gu, 2024) linear-complexity architecture, to make our conclusions generalizable and experiments affordable in computation. To our surprise, this simple scaling study delivers three intriguing discoveries:
@@ 2 Related Work
The landscape of visual backbones underwent another round of significant transformation with the introduction of ViTs (Dosovitskiy et al., 2021) in late 2020, where a novel plain architecture was proposed that treats images akin to language sequences. This model utilizes a simple patchification layer to convert images into sequences of tokens, which are then processed using mechanisms adapted from language models. This approach opened new avenues in handling visual data without the inductive biases inherent in CNNs, demonstrating competitive performance on several benchmarks. The success of ViTs have spurred rapid development and innovations in data-efficient training strategies (Touvron et al., 2021a, 2022a, 2022b), self-supervised learning techniques (Caron et al., 2021; Chen et al., 2021b; Bao et al., 2022; He et al., 2022), vision-language understanding (Radford et al., 2021; Jia et al., 2021; Liu et al., 2023; Alayrac et al., 2022; Yu et al., 2022a), and hierarchical architecture designs (Liu et al., 2021; Chen et al., 2021a; Yuan et al., 2021; Yu et al., 2022b).
Inspired by the patchification design of transformers, there have been many CNN-based (Liu et al., 2022) and State Space Model (Kalman, 1960; Gu et al., 2022, 2021) based architectures (Zhu et al., 2024; Wang et al., 2024a, b) following the same paradigm. Notably, the Mamba (Gu & Dao, 2023; Dao & Gu, 2024) token mixer, due to its advantage of linear complexity, has recently been widely used to explore vision tasks and has achieved competitive results (Zhu et al., 2024; Wang et al., 2024a, b; Ren et al., 2024; Liu et al., 2024; Yang et al., 2024; Hatamizadeh & Kautz, 2024; Huang et al., 2024; Li et al., 2024; Lieber et al., 2024). Among them, the Adventurer (Wang et al., 2024b) architecture, which significantly simplifies the overall model, has demonstrated superior speed compared to the Transformer. In this paper, we employ it as one of the primary experimental models.
Visual architecture scaling. Scaling laws was initially studied in natural language processing (Kaplan et al., 2020). In vision, a similar concept has guided the community to scale up foundational models in both parameter size and data volume. For example, in the age of CNNs, EfficientNets (Tan & Le, 2019, 2021) have proposed to scale-up the models in depth, width, and resolution. These advancements were then integrated into ResNets, leading to nearly Billion-level parameter CNNs (Xie et al., 2020; Kolesnikov et al., 2020; Huang et al., 2019; Bello et al., 2021; Wightman et al., 2021). Scaling the parameter count of Vision Transformers has also shown a great success in modern visual understanding benchmarks and has exhibited state-of-the-art results (Touvron et al., 2021b; Zhou et al., 2021; Zhai et al., 2022; Dehghani et al., 2023).
@@ 3.1 Problem Formulation
To eliminate the influence of different mixer types on the results of patchification scaling, we conduct the main experiments using two visual encoders: the standard ViT (Dosovitskiy et al., 2021) and Adventurer (Wang et al., 2024b). Due to the significant memory and computation challenges posed by the quadratic complexity of self-attention, ViT is only used for context lengths within 4,096 in this work. For longer sequence tasks, we employ Adventurer, a recent Mamba-based (Gu & Dao, 2023; Dao & Gu, 2024) efficient architecture that excels in modeling long range dependencies with linear complexity. Adventurer shares the same plain framework as ViT, with spatial compression only presents in the initial patchification layer, while the key difference is that Adventurer leverages the recent Mamba (Dao & Gu, 2024) module as its token mixer, which has a linear complexity relative to sequence length and allows us to perform pixel tokenization for even the standard 224224 resolution inputs within reasonable computational resources (e.g., 256 A100 GPUs). Remarkably, in our experiments, we form a super-long visual sequence of 50,176 tokens for ImageNet inputs by scaling down the patch size to 11.
@@ 4 Experiments
| Adventurer-Base (Wang et al., 2024b) | 6464 | 69.2 | 77.2 | 80.0 | 80.5 | 80.9 | 4,096 | ||
| Adventurer-Base (Wang et al., 2024b) | 128128 | 79.0 | 81.5 | 81.8 | 82.2 | 82.4 | 16,384 | ||
| Adventurer-Base (Wang et al., 2024b) | 224224 | 82.6 | 83.9 | 84.3 | 84.5 | 84.6 | 50,176 | ||
@@ 4.3 Ablation Studies
Limitations of input size scaling. Compared to scaling down the patch size at a fixed input size, another method—directly scaling up the input size—can achieve a similar effect of reducing the compression rate and extending the token sequence. However, we contend that changing the input size is not a flexible and applicable approach for effective scaling, as its upper bound is easily constrained by the original resolution of the image. For example, in the standard ImageNet classification benchmark, images are resized to 224224 for both training and evaluation stages (Touvron et al., 2021a; Liu et al., 2022; Wang et al., 2024b). This input size has actually compressed the visual information, as the average ImageNet image size is approximately 490430 pixels. Within this range, scaling the input size is generally more effective than scaling the patch size. For example, Adventurer-B/16 with a 448×448 input outperforms Adventurer-B/8 with a 224×224 input by 0.4%, despite having similar parameter counts and FLOPs.
@@ Acknowledgments
We would like to thank TPU Research Cloud (TRC) program and Google Cloud Research Credits program for partially supporting our computing needs.
@@ A. More Technical Details
| Adventurer-Tiny (Wang et al., 2024b) | 256 | 640 | 12 | 12M |
| Adventurer-Small (Wang et al., 2024b) | 512 | 1,280 | 12 | 44M |
| Adventurer-Base (Wang et al., 2024b) | 768 | 1,920 | 12 | 99M |
| Adventurer-Large (Wang et al., 2024b) | 1,024 | 2,560 | 24 | 346M |
| Adventurer-Huge (Wang et al., 2024b) | 1,280 | 3,200 | 32 | 759M |
@@
For Adventurer, we mostly follow its original multi-stage strategy (Wang et al., 2024b) to train our models. Specifically, for 6464 resolution inputs, we simply perform the pretraining stage (shown in Table 8) for 300 epochs for all model sizes. For 128128 resolution inputs, we additionally perform a finetuning stage (shown in Table 9) for enhanced results. For the standard 224224 resolution inputs, we follow the practice of Mamba-Reg (Wang et al., 2024a) and Adventurer (Wang et al., 2024b) to load the pretrained model at 128128, performing an intermediate training stage (shown in Table 10) for 100 epochs and then a finetuning stage for 20 epochs. We highlight that this multi-stage training strategy is highly efficient for our experiments as we can fully exploit the models pretrained at lower resolutions. For example, we only need to train the 224224-input models for 120 epochs since we can load the weights pretrained at 128128 resolution inputs.

### v2

@@ 1 Introduction
Thanks to the rapid advancements in hardware, efficient attention mechanisms (Dao et al., 2022; Kwon et al., 2023), as well as linear-complexity structures (Katharopoulos et al., 2020; Peng et al., 2023; Gu & Dao, 2023), we can now extensively validate the impact of patchification at a standard input size (e.g., 224224 for ImageNet) with manageable computing resources. We conduct a series of straightforward scaling experiments on patchification, gradually reducing the model’s patch size from the typical 1616 down to 11 to lower the compression rate and observe how performance changes. We employ both ViT and Adventurer (Wang et al., 2024c), a Mamba-based (Dao & Gu, 2024) linear-complexity architecture, to make our conclusions generalizable and experiments affordable in computation. To our surprise, this simple scaling study delivers three intriguing discoveries:
@@ 2 Related Work
The landscape of visual backbones underwent another round of significant transformation with the introduction of ViTs (Dosovitskiy et al., 2021) in late 2020, where a novel plain architecture was proposed that treats images akin to language sequences. This model utilizes a simple patchification layer to convert images into sequences of tokens, which are then processed using mechanisms adapted from language models. This approach opened new avenues in handling visual data without the inductive biases inherent in CNNs, demonstrating competitive performance on several benchmarks. The success of ViTs have spurred rapid development and innovations in data-efficient training strategies (Touvron et al., 2021a, 2022a, 2022b), self-supervised learning techniques (Caron et al., 2021; Chen et al., 2021b; Bao et al., 2022; He et al., 2022; Ren et al., 2024b), vision-language understanding (Radford et al., 2021; Jia et al., 2021; Liu et al., 2023; Alayrac et al., 2022; Yu et al., 2022a; Wang et al., 2024a), and hierarchical architecture designs (Liu et al., 2021; Chen et al., 2021a; Yuan et al., 2021; Yu et al., 2022b).
Inspired by the patchification design of transformers, there have been many CNN-based (Liu et al., 2022) and State Space Model (Kalman, 1960; Gu et al., 2022, 2021) based architectures (Zhu et al., 2024; Wang et al., 2024b, c) following the same paradigm. Notably, the Mamba (Gu & Dao, 2023; Dao & Gu, 2024) token mixer, due to its advantage of linear complexity, has recently been widely used to explore vision tasks and has achieved competitive results (Zhu et al., 2024; Wang et al., 2024b, c; Ren et al., 2024a; Liu et al., 2024; Yang et al., 2024; Hatamizadeh & Kautz, 2024; Huang et al., 2024; Li et al., 2024; Ren et al., 2024c; Lieber et al., 2024; Wei & Chellappa, 2025). Among them, the Adventurer (Wang et al., 2024c) architecture, which significantly simplifies the overall model, has demonstrated superior speed compared to the Transformer. In this paper, we employ it as one of the primary experimental models.
Visual architecture scaling. Scaling laws was initially studied in natural language processing (Kaplan et al., 2020). In vision, a similar concept has guided the community to scale up foundational models in both parameter size and data volume. For example, in the age of CNNs, EfficientNets (Tan & Le, 2019, 2021) have proposed to scale-up the models in depth, width, and resolution. These advancements were then integrated into ResNets, leading to nearly Billion-level parameter CNNs (Xie et al., 2020; Kolesnikov et al., 2020; Huang et al., 2019; Bello et al., 2021; Wightman et al., 2021). Scaling the parameter count of Vision Transformers has also shown a great success in modern visual understanding benchmarks and has exhibited state-of-the-art results (Touvron et al., 2021b; Zhou et al., 2021; Zhai et al., 2022; Dehghani et al., 2023). More recently, Nguyen et al. introduce Pixel Transformers, with standard patchification grids scaled down to pixels, showcasing promising scaling results for low-resolution (e.g., 3232) input images.
@@ 3.1 Problem Formulation
To eliminate the influence of different mixer types on the results of patchification scaling, we conduct the main experiments using two visual encoders: the standard ViT (Dosovitskiy et al., 2021) and Adventurer (Wang et al., 2024c). Due to the significant memory and computation challenges posed by the quadratic complexity of self-attention, ViT is only used for context lengths within 4,096 in this work. For longer sequence tasks, we employ Adventurer, a recent Mamba-based (Gu & Dao, 2023; Dao & Gu, 2024) efficient architecture that excels in modeling long range dependencies with linear complexity. Adventurer shares the same plain framework as ViT, with spatial compression only presents in the initial patchification layer, while the key difference is that Adventurer leverages the recent Mamba (Dao & Gu, 2024) module as its token mixer, which has a linear complexity relative to sequence length and allows us to perform pixel tokenization for even the standard 224224 resolution inputs within reasonable computational resources (e.g., 256 A100 GPUs). Remarkably, in our experiments, we form a super-long visual sequence of 50,176 tokens for ImageNet inputs by scaling down the patch size to 11.
@@ 4 Experiments
| Adventurer-Base (Wang et al., 2024c) | 6464 | 69.2 | 77.2 | 80.0 | 80.5 | 80.9 | 4,096 | ||
| Adventurer-Base (Wang et al., 2024c) | 128128 | 79.0 | 81.5 | 81.8 | 82.2 | 82.4 | 16,384 | ||
| Adventurer-Base (Wang et al., 2024c) | 224224 | 82.6 | 83.9 | 84.3 | 84.5 | 84.6 | 50,176 | ||
@@ 4.3 Ablation Studies
Limitations of input size scaling. Compared to scaling down the patch size at a fixed input size, another method—directly scaling up the input size—can achieve a similar effect of reducing the compression rate and extending the token sequence. However, we contend that changing the input size is not a flexible and applicable approach for effective scaling, as its upper bound is easily constrained by the original resolution of the image. For example, in the standard ImageNet classification benchmark, images are resized to 224224 for both training and evaluation stages (Touvron et al., 2021a; Liu et al., 2022; Wang et al., 2024c). This input size has actually compressed the visual information, as the average ImageNet image size is approximately 490430 pixels. Within this range, scaling the input size is generally more effective than scaling the patch size. For example, Adventurer-B/16 with a 448×448 input outperforms Adventurer-B/8 with a 224×224 input by 0.4%, despite having similar parameter counts and FLOPs.
@@ Acknowledgments
This work is supported by ONR: N00014-21-1-2690 and National Eye Institute (NEI) with Award ID: R01EY037193. We would like to thank TPU Research Cloud (TRC) program and Google Cloud Research Credits program for partially supporting our computing needs.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. Our experiments involved approximately 50,000 A100 GPU hours, which is considered a modest level of resource consumption compared to large-scale vision or language model research. While there are many potential societal consequences of our work, none are significant enough to warrant specific highlighting in this context. We believe the ethical impacts and societal implications are well-aligned with the advancement of machine learning technology.
@@ A. More Technical Details
| Adventurer-Tiny (Wang et al., 2024c) | 256 | 640 | 12 | 12M |
| Adventurer-Small (Wang et al., 2024c) | 512 | 1,280 | 12 | 44M |
| Adventurer-Base (Wang et al., 2024c) | 768 | 1,920 | 12 | 99M |
| Adventurer-Large (Wang et al., 2024c) | 1,024 | 2,560 | 24 | 346M |
| Adventurer-Huge (Wang et al., 2024c) | 1,280 | 3,200 | 32 | 759M |
@@
For Adventurer, we mostly follow its original multi-stage strategy (Wang et al., 2024c) to train our models. Specifically, for 6464 resolution inputs, we simply perform the pretraining stage (shown in Table 8) for 300 epochs for all model sizes. For 128128 resolution inputs, we additionally perform a finetuning stage (shown in Table 9) for enhanced results. For the standard 224224 resolution inputs, we follow the practice of Mamba-Reg (Wang et al., 2024b) and Adventurer (Wang et al., 2024c) to load the pretrained model at 128128, performing an intermediate training stage (shown in Table 10) for 100 epochs and then a finetuning stage for 20 epochs. We highlight that this multi-stage training strategy is highly efficient for our experiments as we can fully exploit the models pretrained at lower resolutions. For example, we only need to train the 224224-input models for 120 epochs since we can load the weights pretrained at 128128 resolution inputs.